OpenStudy (anonymous):

x(x-2)(x+3)=0 plz help!!!

OpenStudy (anonymous):

What about it? Are you trying to solve for x?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Use zero-product-property: set each factor =0, and solve individually.

OpenStudy (anonymous):

Hmmm how can I do that?

OpenStudy (anonymous):

If you don't know how to do that, then you shouldn't be attempting this problem. (no offense)

OpenStudy (anonymous):

Well I get it a little bit

OpenStudy (anonymous):

Basically, you have to understand that this means that you have to find the value(s) of x that make the equation equal to zero. To do this, you have to understand that 0 multiplied by anything is still 0. Therefore, to get x(x - 2)(x + 3) to equal to zero, either x, x - 2, or x + 3 have to equal 0. For instance, x = 0 is a solution. Since 0 times anything is 0, you can use 0. 0(0 - 2)(0 + 3) = 0(-2)(3) = 0. You can even leave it as 0(x - 2)(x - 3) = 0 because 0 times anything is still 0. Now set x - 2 equal to 0 and x + 3 equal to 0. x = 0 x - 2 = 0 x = 2 x + 3 = 0 x = -3 Therefore, x = 0, 2, and -3.

OpenStudy (anonymous):

Ok yea I get it thank you soooo much!

OpenStudy (anonymous):

You're welcome!